OpenStudy (anonymous):

Use De Moivre's Theroem to find the square root of -5+12i.

OpenStudy (stacey):

You want to rewrite that in the form \[r(\cos x+i\sin x)\]

OpenStudy (anonymous):

Yes. But I could not find the arg(-5+12i). Could you please show me how to?

OpenStudy (stacey):

\[r=\sqrt{(-5)^{2}+12^{2}}\]

OpenStudy (anonymous):

Ok. So r=13, what next?

OpenStudy (stacey):

If we factor 13 out then we have \[13(\frac{ -5 }{ 13 }+\frac{ 12i }{13 })\]and so we want to find at what angle cosine equals -5/13 and sine equals 12/13.

OpenStudy (anonymous):

Get it. How do we find that? :O

OpenStudy (stacey):

A calculator. Enter\[\cos^{-1} (-5/13)\]

OpenStudy (stacey):

and\[\sin^{-1} (12/13)\]

OpenStudy (anonymous):

112.62 and 67.38

OpenStudy (stacey):

Both should have the same sign, so we need to adjust the angles accordingly

OpenStudy (anonymous):

The book gives an answer in this form. That was why I got confused. I don't get what this 0.98^c stands for? :O \[\pm \sqrt{5}\left\{ \cos(0.98^{c})+isin(0.98^{c}) \right\}\]

OpenStudy (anonymous):

Sorry. That's \[\sqrt{13}\]

OpenStudy (stacey):

That does not make sense. I wonder if there is some sort of typing error.

OpenStudy (stacey):

The degrees should be approximately 112.62 for both sin and cos.

OpenStudy (anonymous):

No. There are other questions like this and answers in this format. I'm just stuck. :-(

OpenStudy (stacey):

Before you divide by 2, that is.

OpenStudy (stacey):

Do they all have c's instead of degree signs?

OpenStudy (anonymous):

Np. Some answers are in the form \[\pm \sqrt{12}(\cos \pi/12+isin \pi/12)\]

OpenStudy (stacey):

Okay. Your book is working with radians instead of degrees.

OpenStudy (stacey):

You should change the mode of your calculator to do radians.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

I got1.176 for sin^-1 (12/13) and 1.965 for cos^-1(-5/13).

OpenStudy (stacey):

That makes sense, but only one of those is correct.